OpenStudy (justtryingtosurvive):

WILL MEDAL AND FAN NEED HELP DESPERATELY A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 10(1.02)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.04 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5, and what does it represent?

OpenStudy (wanell):

PART A) 11.04 = 10(1.02)^n 1.104 = 1.02^n ln 1.104 = ln 1.02^n ln 1.104 = n ln 1.02 n = ln 1.104/ ln 1.02 n = 4.99630409516 4.99 can be rounded to 5. So a reasonable domain would be 0 ≤ x < 5 PART B) f(0) = 10(1.02)^0 f(0) = 10(1) f(0) = 10 The y-intercept represents the height of the plant when they began the experiment. PART C) f(1) = 10(1.02)^1 f(1) = 10(1.02) f(1) = 10.2 (1, 10.2) f(5) = 10(1.02)^5 f(5) = 10(1.1040808) f(5) = 11.040808 (5, 11.040808) Find the slope between those two points.

OpenStudy (wanell):

The slope represents how much the plant is growing every day.